OpenStudy (highschoolmom2010):

Find the volume in terms of pi of each sphere with the given surface area. 36 pi in ^2

OpenStudy (highschoolmom2010):

\[SA=4 \pi r^2=36 \pi (inches) ^2\]

OpenStudy (tkhunny):

You can just use the Surface Area to solve for 'r', but I would like to suggest a somewhat off-handed approach that might inspire you toward creativity. Volume = \(\dfrac{4}{3}\pi r^{3} = 4\pi r^{2}\cdot \dfrac{r}{3} = SA\cdot\dfrac{r}{3}\) You must find r/3. If \(4\pi r^{2} = 36\pi\;then\;\pi r^{2} = 9\pi\;and\;r^{2} = 9 \implies r = 3\;and\;\dfrac{r}{3} = 1\)

OpenStudy (highschoolmom2010):

ok so how am i supposed to find the surface area using ^^

OpenStudy (tkhunny):

?? You are GIVEN the surface area. It's in the problem statement.

OpenStudy (highschoolmom2010):

oh